# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
1:Input: Offline dataset $\mathcal{D}_{\rm off}$, Online environment $\mathcal{M}_{\rm on}$, max interaction steps $T_{\rm max}$, target model training frequency $\mathsf{train\_freq}$, source data selection ratio $\xi$, gap reward scale $\beta$, batch size $B$, behavior cloning weight $\omega$, warmup steps $\mathsf{warmup\_steps}$, learning rate $\mathsf{lr}$, target update rate $\varpi$
2:Initialization: Policy $\pi_{\varphi}$, Q-functions $\{Q_{\varsigma_{i}}\}_{i=1,2}$, target Q-functions $\{Q_{\varsigma_{i}}^{\rm tgt}\}_{i=1,2}\leftarrow\{Q_{\varsigma_{i}}\}_{i=1,2}$, target replay buffer $\mathcal{D}_{\rm on}$
3:Pretrain offline flow model $\psi_{\theta}(x,t\mid s,a)$ on $\mathcal{D}_{\rm off}$ via Algorithm 2
4:for $t=1$ to $T_{\rm max}$ do
5:  Sample transition $(s_{\rm on},a_{\rm on},r_{\rm on},s^{\prime}_{\rm on})\sim\mathcal{M}_{\rm on}$ using policy $\pi_{\varphi}$
6:  Update replay buffer: $\mathcal{D}_{\rm on}\leftarrow\mathcal{D}_{\rm on}\cup\{(s_{\rm on},a_{\rm on},r_{\rm on},s^{\prime}_{\rm on})\}$
7:  if $t\bmod\mathsf{train\_freq}=0$ and $t>\mathsf{warmup\_steps}$ then
8:   Train target flow $\psi_{\phi}(x,t\mid s,a)$ on $\mathcal{D}_{\rm on}$ via Algorithm 3
9:  end if
10:  for $k=1$ to $K$ do
11:   Sample $b_{\rm off}=\{(s_{\rm off}^{(i)},a_{\rm off}^{(i)},r_{\rm off}^{(i)},s_{\rm off}^{\prime(i)})\}_{i=1}^{B}\sim D_{\rm off}$
12:   Sample $b_{\rm on}=\{(s_{\rm on}^{(i)},a_{\rm on}^{(i)},r_{\rm on}^{(i)},s_{\rm on}^{\prime(i)})\}_{i=1}^{B}\sim\mathcal{D}_{\rm on}$
13:   Estimate $\widehat{\Delta}(s_{\rm off}^{(i)},a_{\rm off}^{(i)})$ for each $(s_{\rm off}^{(i)},a_{\rm off}^{(i)})\in b_{\rm off}$ via Eq. 3
14:   Select top $\xi\%$ of $b_{\rm off}$ with lowest gap: $\tilde{b}_{\rm off}$
15:   For each $(s,a,s^{\prime},r)\in\tilde{b}_{\rm off}$, update reward: $r\leftarrow r+\beta\widehat{\Delta}(s,a)$
16:   for $i=1,2$ do
17:     Compute Bellman target using target Q-networks:
$y=r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\varphi}(\cdot|s^{\prime})}\left[\min_{j=1,2}Q^{\rm tgt}_{\varsigma_{j}}(s^{\prime},a^{\prime})-\alpha\log\pi_{\varphi}(a^{\prime}|s^{\prime})\right]$
18:     Compute Q-function loss gradient:
$\nabla_{\varsigma_{i}}\mathcal{L}_{Q}\leftarrow$ $\frac{1}{B}\sum_{(s^{\rm on},a_{\rm on},s^{\prime}_{\rm on},r_{\rm on})\in b_{\rm on}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s_{\rm on},a_{\rm on})-y\right)^{2}$
$+\frac{1}{|\tilde{b}_{\rm off}|}\sum_{(s_{\rm off},a_{\rm off},s^{\prime}_{\rm off},r_{\rm off})\in\tilde{b}_{\rm off}}\nabla_{\varsigma_{i}}\left(Q_{\varsigma_{i}}(s_{\rm off},a_{\rm off})-y\right)^{2}$
19:     Update Q-network: $\varsigma_{i}\leftarrow\varsigma_{i}-\mathsf{lr}\nabla_{\varsigma_{i}}\mathcal{L}_{Q}$
@@
21:   # Soft update of target Q-networks
@@
23:     $\varsigma_{i}^{\rm tgt}\leftarrow\varpi\varsigma_{i}+(1-\varpi)\varsigma_{i}^{\rm tgt}$
@@
25:   Compute BC weight:
[…]
26:   Compute policy loss gradient:
[…]
$+\frac{1}{|b_{\rm off}|}\sum_{(s,a)\in b_{\rm off}}\nabla_{\varphi}\left\|a-\bar{a}\right\|^{2}$
27:   where $\tilde{a}\sim\pi_{\varphi}(\cdot|\tilde{s}),\quad\bar{a}\sim\pi_{\varphi}(\cdot|s)$ are independent samples.
28:   Update policy: $\varphi\leftarrow\varphi+\mathsf{lr}\nabla_{\varphi}\mathcal{L}_{\pi}$
29:  end for
30:end for
31:Output: Learned policy $\pi_{\varphi}$
@@
Proof of Lemma  2.2 is already provided as an intermediate step for Theorem A.4 in [lyu2024cross], the offline performance bound case. We include their proof here with slight modifications for completeness.
We first cite the following two necessary lemmas also used in [lyu2024cross]. ∎
@@
Proof. This is Lemma C.2 in [xu2023cross]. $\square$
@@
Proof. This is Lemma B.3 in [lyu2024cross]. $\square$
@@
One can analogously replace $W_{2}(p_{0},p_{\rm on})$ by $D_{2}(p_{0},p_{\rm on})$ when bounding 14, using the standard inequality trick $\sqrt{x}\leq\tfrac{1}{2}(x+1)$ for all $x\geq 0$. We use $W_{2}$ metric because it’s more commonly adopted in the literature, but $D_{2}$ leads to a tighter bound since we no longer needs the constant $C_{\rm TR}$. We do not expand the $D_{2}$ case here, as the argument mirrors the $W_{2}$ case.
@@
For any fixed policy $\pi$, we cite Theorem 3.1 of [neufeld2023bounding] for the following bound. Under R1R3, we have
@@
We provide the formal version of Proposition 3.3 as below.
Proposition I.1 (Shared-Latent Coupling Approximates $W_{2}$-Optimal Transport).
We first introduce some notations for simplicity:
Notation. Fix a state–action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, and omit $(s,a)$ in the following for notational simplicity. Let the latent distribution be $p_{0}(x_{0})=\mathcal{N}(0,\mathbf{I})$. Define the offline transition distribution as $p_{1}(x_{1}):=p_{\rm off}(s^{\prime}|s,a)$, and the online transition distribution as $p_{2}(x_{2}):=p_{\rm{on}}(s^{\prime}|s,a)$.
We denote the pretrained source flow $\psi_{\theta}(x,t|s,a)$ by $\psi_{\theta}(x,t)$, and similarly write the target flow $\psi_{\phi}(x,t|s,a)$ as $\psi_{\phi}(x,t)$. Let $\psi_{\phi}^{k}(x,t)$ denote the target flow model trained on a batch of size $k$ from the target domain.
Assumptions.
(A1) The distributions $p_{1}$ and $p_{2}$ have bounded support: there exists $C>0$ such that $\|x\|\leq C$ for all $x\in\mathrm{supp}(p_{1})\cup\mathrm{supp}(p_{2})$.
(A2) The distribution $p_{1}$ admits a density, and the optimal transport map between $p_{1}$ and $p_{2}$ under the quadratic cost is continuous.
(A3) At each training iteration, we compute the optimal transport plan for the current minibatch.
Result. Let $W_{2}^{2}(p_{1},p_{2})$ denote the squared 2-Wasserstein distance between $p_{1}$ and $p_{2}$, defined as
$W_{2}^{2}(p_{1},p_{2}):=\min_{q\in\Pi(p_{1},p_{2})}\mathbb{E}_{(x_{1},x_{2})\sim q}\left[\|x_{2}-x_{1}\|_{2}^{2}\right],$
where $\Pi(p_{1},p_{2})$ denotes the set of all couplings (i.e., joint distributions) with marginals $p_{1}$ and $p_{2}$.
Then, in the limit as $\eta\to\infty$ (in the cost function of Algorithm 3) and $k\to\infty$, we have
$\lim_{k\to\infty}\mathbb{E}_{x_{0}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\psi_{\theta}(x_{0},1)-\psi^{k}_{\phi}\left(\psi_{\theta}(x_{0},1),2\right)\right\|^{2}\right]=W_{2}^{2}(p_{1},p_{2}),$
where $\psi_{\theta}$ denotes the source flow and $\psi_{\phi}^{k}$ is the target flow trained on a batch of size $k$.

### v4

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
1:Input: Offline dataset $\mathcal{D}_{\rm off}$, online environment $\mathcal{M}_{\rm on}$, max interaction steps $T_{\max}$, update frequency $\mathsf{train\_freq}$, offline selection ratio $\xi\in(0,1]$, gap reward scale $\beta$, batch size $B$, behavior cloning weight $\omega$, warmup steps $\mathsf{warmup\_steps}$, gradient steps per interaction $K$, learning rate $\mathsf{lr}$, target update rate $\varpi$, discount $\gamma$, entropy temperature $\alpha$.
2:Initialize: Policy $\pi_{\varphi}$, critics $\{Q_{\varsigma_{i}}\}_{i=1,2}$, target critics $\{Q^{\rm tgt}_{\varsigma_{i}}\}_{i=1,2}\leftarrow\{Q_{\varsigma_{i}}\}_{i=1,2}$, online replay buffer $\mathcal{D}_{\rm on}\leftarrow\emptyset$.
3:Pretrain offline flow $\psi_{\theta}(x,t\mid s,a)$ on $\mathcal{D}_{\rm off}$ via Algorithm 2.
4:for $t=1$ to $T_{\max}$ do
5:  Interact with $\mathcal{M}_{\rm on}$ using $\pi_{\varphi}$ and store transition $(s,a,r,s^{\prime})$ into $\mathcal{D}_{\rm on}$. $\triangleright$ $a\sim\pi_{\varphi}(\cdot\mid s)$, $(r,s^{\prime})\sim\mathcal{M}_{\rm on}(s,a)$
6:  if $t>\mathsf{warmup\_steps}$ and $t\bmod\mathsf{train\_freq}=0$ then
7:   Train online flow $\psi_{\phi}(x,t\mid s,a)$ on $\mathcal{D}_{\rm on}$ via Algorithm 3.
8:  end if
9:  for $k=1$ to $K$ do
10:   Sample offline minibatch $b_{\rm off}=\{(s^{(i)}_{\rm off},a^{(i)}_{\rm off},r^{(i)}_{\rm off},s_{\rm off}^{\prime(i)})\}_{i=1}^{B}\sim\mathcal{D}_{\rm off}$.
11:   Sample online minibatch $b_{\rm on}=\{(s^{(i)}_{\rm on},a^{(i)}_{\rm on},r^{(i)}_{\rm on},s_{\rm on}^{\prime(i)})\}_{i=1}^{B}\sim\mathcal{D}_{\rm on}$.
12:   Estimate dynamics gap $\widehat{\Delta}(s^{(i)}_{\rm off},a^{(i)}_{\rm off})$ for each $(s^{(i)}_{\rm off},a^{(i)}_{\rm off})$ via Eq. (4).
13:   Select the lowest-gap subset $\tilde{b}_{\rm off}\subset b_{\rm off}$ with $|\tilde{b}_{\rm off}|=\lceil\xi B\rceil$.
14:   Shape rewards for selected offline transitions: for each $(s,a,r,s^{\prime})\in\tilde{b}_{\rm off}$ set $\tilde{r}\leftarrow r+\beta\,\widehat{\Delta}(s,a)$.
15:   # Critic update (use online batch + selected offline batch)
16:   Define critic batch $b_{Q}\leftarrow b_{\rm on}\cup\tilde{b}_{\rm off}$ with rewards $r$ for online and $\tilde{r}$ for selected offline.
17:   for $i=1,2$ do
18:     Compute Bellman targets for $(s,a,r,s^{\prime})\in b_{Q}$:
$y(s,a,r,s^{\prime})\;=\;r+\gamma\,\mathbb{E}_{a^{\prime}\sim\pi_{\varphi}(\cdot\mid s^{\prime})}\Big[\min_{j=1,2}Q^{\rm tgt}_{\varsigma_{j}}(s^{\prime},a^{\prime})-\alpha\log\pi_{\varphi}(a^{\prime}\mid s^{\prime})\Big].$
19:     Update critic by gradient descent:
$\varsigma_{i}\leftarrow\varsigma_{i}-\mathsf{lr}\,\nabla_{\varsigma_{i}}\Bigg(\frac{1}{|b_{Q}|}\sum_{(s,a,r,s^{\prime})\in b_{Q}}\big(Q_{\varsigma_{i}}(s,a)-y(s,a,r,s^{\prime})\big)^{2}\Bigg).$
@@
21:   # Soft update target critics
@@
23:     $\varsigma_{i}^{\rm tgt}\leftarrow\varpi\,\varsigma_{i}+(1-\varpi)\,\varsigma_{i}^{\rm tgt}$.
@@
25:   # Actor update (policy improvement + behavior cloning)
26:   Define actor state batch $b_{\pi}\leftarrow\{s:(s,\cdot,\cdot,\cdot)\in b_{\rm on}\cup\tilde{b}_{\rm off}\}$.
27:   Sample $\tilde{a}\sim\pi_{\varphi}(\cdot\mid s)$ for each $s\in b_{\pi}$.
28:   Set normalization weight
$\lambda\;=\;\frac{\omega}{\frac{1}{|b_{\pi}|}\sum_{s\in b_{\pi}}\left|\min_{i=1,2}Q_{\varsigma_{i}}(s,\tilde{a})\right|+\epsilon},\qquad\epsilon>0.$
29:   Minimize actor loss:
$\mathcal{L}_{\pi}(\varphi)\;=\;$ $\ \frac{1}{|b_{\pi}|}\sum_{s\in b_{\pi}}\Big(\alpha\log\pi_{\varphi}(\tilde{a}\mid s)-\min_{i=1,2}Q_{\varsigma_{i}}(s,\tilde{a})\Big)\;+\;\lambda\,\frac{1}{|b_{\rm off}|}\sum_{(s,a,\cdot,\cdot)\in b_{\rm off}}\|\;a-\bar{a}\;\|_{2}^{2},$
30:   where $\tilde{a}\sim\pi_{\varphi}(\cdot\mid s)$ for $s\in b_{\pi}$, and $\bar{a}\sim\pi_{\varphi}(\cdot\mid s)$ for $(s,a)\in b_{\rm off}$.
31:   $\varphi\leftarrow\varphi-\mathsf{lr}\,\nabla_{\varphi}\mathcal{L}_{\pi}(\varphi)$.
32:  end for
33:end for
34:Output: Learned policy $\pi_{\varphi}$.
@@
Proof of Lemma  2.2 is already provided as an intermediate step for Theorem A.4 in [37], the offline performance bound case. We include their proof here with slight modifications for completeness.
We first cite the following two necessary lemmas also used in [37]. ∎
@@
Proof. This is Lemma C.2 in [65]. $\square$
@@
Proof. This is Lemma B.3 in [37]. $\square$
@@
One can analogously replace $W_{2}(p_{0},p_{\rm on})$ by $D_{2}(p_{0},p_{\rm on})$ when bounding 15, using the standard inequality trick $\sqrt{x}\leq\tfrac{1}{2}(x+1)$ for all $x\geq 0$. We use $W_{2}$ metric because it’s more commonly adopted in the literature, but $D_{2}$ leads to a tighter bound since we no longer needs the constant $C_{\rm TR}$. We do not expand the $D_{2}$ case here, as the argument mirrors the $W_{2}$ case.
@@
For any fixed policy $\pi$, we cite Theorem 3.1 of [45] for the following bound. Under R1R3, we have
@@
We provide a rigorous justification for Proposition 3.3. The key technical point is that the augmented optimal transport (OT) problem on triples $(s,a,s^{\prime})$ with cost
$c_{\eta}\big((s,a,x),(s^{\prime},a^{\prime},x^{\prime})\big)=\|x-x^{\prime}\|_{2}^{2}+\eta\big(\|s-s^{\prime}\|_{2}^{2}+\|a-a^{\prime}\|_{2}^{2}\big)$
reduces, as $\eta\to\infty$, to conditional quadratic OT on the next-state variable $x=s^{\prime}$ given each fixed $(s,a)$, provided that the $(s,a)$-marginals match.
I.1 Formal Proposition
Proposition I.1 (Shared-latent coupling recovers conditional $W_{2}$).
Setup. Let $\mathcal{Y}:=\mathcal{S}\times\mathcal{A}$ denote the conditioning space and $\mathcal{X}:=\mathcal{S}$ denote the next-state space. Let $\mu$ be a probability measure on $\mathcal{Y}$. Let $\alpha,\beta$ be probability measures on $\mathcal{Y}\times\mathcal{X}$ such that their $\mathcal{Y}$-marginals coincide:
$\pi_{sa}\#\alpha=\pi_{sa}\#\beta=\mu,$ (42)
where $\pi_{sa}(y,x)=y$. Assume $\alpha$ and $\beta$ admit disintegrations with respect to $\mu$:
$\alpha(dy,dx)=\mu(dy)\,\alpha_{y}(dx),\qquad\beta(dy,dx)=\mu(dy)\,\beta_{y}(dx),$ (43)
where $y=(s,a)\in\mathcal{Y}$ and $x\in\mathcal{X}$.
Augmented OT problem. For $\eta>0$, define the augmented quadratic cost on $(\mathcal{Y}\times\mathcal{X})^{2}$ by
$c_{\eta}\big((y,x),(y^{\prime},x^{\prime})\big):=\|x-x^{\prime}\|_{2}^{2}+\eta\|y-y^{\prime}\|_{2}^{2},$ (44)
where $\|y-y^{\prime}\|_{2}^{2}:=\|s-s^{\prime}\|_{2}^{2}+\|a-a^{\prime}\|_{2}^{2}$. Define the corresponding OT value
$W_{2,\eta}^{2}(\alpha,\beta):=\inf_{\gamma\in\Pi(\alpha,\beta)}\int c_{\eta}\big((y,x),(y^{\prime},x^{\prime})\big)\,d\gamma,$ (45)
where $\Pi(\alpha,\beta)$ is the set of couplings on $(\mathcal{Y}\times\mathcal{X})^{2}$ with marginals $\alpha$ and $\beta$.
Assumptions (for $\eta\to\infty$ reduction). Assume $\alpha$ and $\beta$ have finite second moments, that $W_{2}(\alpha_{y},\beta_{y})<\infty$ for $\mu$-a.e. $y$, and that the conditional squared Wasserstein cost is $\mu$-integrable:
$\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy)<\infty.$ (46)
Claim 1 (conditional decomposition). As $\eta\to\infty$,
$\lim_{\eta\to\infty}W_{2,\eta}^{2}(\alpha,\beta)=\int_{\mathcal{Y}}W_{2}^{2}(\alpha_{y},\beta_{y})\,\mu(dy).$ (47)
Claim 2 (structure of limiting couplings). Let $\gamma_{\eta}\in\arg\min_{\gamma\in\Pi(\alpha,\beta)}\int c_{\eta}\,d\gamma$ be any sequence of optimal couplings. Then there exists a subsequence $\gamma_{\eta_{n}}\Rightarrow\gamma_{\infty}$ weakly such that:
  1. (i)
    $\gamma_{\infty}$ is supported on the diagonal set $\{(y,x;y,x^{\prime}):\ y=y^{\prime}\}$;
  2. (ii)
    $\gamma_{\infty}$ admits a disintegration
    $\gamma_{\infty}(dy,dx,dx^{\prime})=\mu(dy)\,\gamma_{y}(dx,dx^{\prime}),$
    where $\gamma_{y}\in\Pi(\alpha_{y},\beta_{y})$ is $W_{2}$-optimal for $\mu$-a.e. $y$.
OT-FM implication (per fixed conditioning $y$). Fix any $y=(s,a)$ and let $p_{0}=\mathcal{N}(0,I)$. Assume the offline flow satisfies
$\psi^{\rm off}_{\theta}(\cdot,1\mid y)\#p_{0}=\alpha_{y}.$
For $\eta>0$ and $k\in\mathbb{N}$, let $\psi^{{\rm on},k,\eta}_{\phi}(\cdot,2\mid y)$ denote the online flow obtained by running OT-FM to convergence with minibatch size $k$, where each iteration computes an optimal transport plan between two empirical measures of size $k$ under the augmented cost (44).
Assume the standard OT-FM regularity conditions hold for this fixed $y$ (e.g., those in Theorem 4.2 of [47]), including:
  1. (A1)
    $\alpha_{y}$ and $\beta_{y}$ have bounded support;
  2. (A2)
    $\alpha_{y}$ admits a density and the quadratic OT map from $\alpha_{y}$ to $\beta_{y}$ exists and is continuous;
  3. (A3)
    the minibatch OT plan is solved exactly at each iteration.
Then, for $\mu$-a.e. $y$, we have the two-limit convergence
$\lim_{\eta\to\infty}\;\lim_{k\to\infty}\;\mathbb{E}_{x_{0}\sim p_{0}}\Big[\big\|\psi^{\rm off}_{\theta}(x_{0},1\mid y)-\psi^{{\rm on},k,\eta}_{\phi}\!\big(\psi^{\rm off}_{\theta}(x_{0},1\mid y),2\mid y\big)\big\|_{2}^{2}\Big]=W_{2}^{2}(\alpha_{y},\beta_{y}).$ (48)
I.2 Proof